if x=2+2^(2/3)+2^(1/3)
then x^3-6x^2+6x= ?
Please show the steps

Grade:11

4 Answers

Aman Bansal
12 years ago

dear Raunak,

the answer to your question is 2

we have x-2=22/3-21/3

then cubing both sides to get,

(x-2)3= 6(x-2)

thus expnding to get,

x3-6x2+6x=2

BEST OF LUCK

Please approve the answer...!!!!!!
We are all IITians and here to help you in your IIT JEE preparation.
Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..  
AMAN BANSAL

Aman Bansal
12 years ago

dear raunak,

I have already posted reply to your question.

BEST OF LUCK

Please approve the answer...!!!!!!
We are all IITians and here to help you in your IIT JEE preparation.
Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..  
AMAN BANSAL

sona
6 years ago

Given that x = 2 + 2^(2/3) + 2^(1/3)

so x - 2 = 2^(2/3) + 2^(1/3)

=(x-2)^3 = [2^(2/3)+2^(1/3)]^3

using (a-b)^3= a^3 - b^3 - 3ab(a-b)

and (a+b)^3= a^3 + b^3 + 3ab(a+b) formulae.

or x^3 - 8 - 6x(x-2) = 2^2 + 2^1 + 3*[2^{(2/3)+(1/3)}[2^(2/3)+2^(1/3)

or x^3 - 8 - 6x^2 + 12x = 4 + 2 + 6(x-2)

or x^3 - 8 -6x^2 + 12x = 6 + 6x - 12

or x^3 - 6x^2 +6x = 2

Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your problem.
 
We have, x = 2 + 22/3 + 21/3
or, x – 2 = 22/3 + 21/3
cubing both sides, we get,
x3 – 8 – 6x(x – 2) = 22 + 2 + 3x23/3(22/3 + 21/3)
x3 – 6x2 + 12x – 8 = 6 + 6(x – 2)                     ( since, 22/3 + 21/3 = x – 2)
hence, x3 – 6x2 + 6x = 6 + 8 -12 = 2
 
Hope it helps.
Thanks and regards,
Kushagra